OpenStudy (anonymous):

How many cats does a woman have when she gives the answer seven-eighths of my cats plus four. I can't determine how to work this problem

OpenStudy (lgbasallote):

i think this question is missing some info. can you check?

OpenStudy (anonymous):

This is actually the question...I It reads: A superhero asks his nemesis how many cats she has. She answers with a riddle "Seven-eighths of my cats plus four." How many cats does the nemesis have?

OpenStudy (lgbasallote):

let x be the number of her cats \[x = \frac 78 x + 4\] subtract 7/8 x from both sides \[x - \frac 78 x = 4\] change to common denominators \[\implies \frac 88 x - \frac 78 x = 4\] combine.. \[\implies \frac 18 x = 4\] cross multiply \[\implies x = 4 \times 8\] does that help?

OpenStudy (anonymous):

yes thank you